G2p Target Minutes

From Hall A Wiki
Jump to: navigation, search

For the target meetings prior to 04/07/2011 see here

Minutes of the weekly analysis meetings

11/03/2011: Attendees: C.Keith, K.Allada, J.Zhang, M.Huang, P. Zhu, C. Gu, M. Cummings, A. Camsonne, D. Higinbotham, J. Pierce et. al
On Phone: J.Maxwell, D. Crabb

JLab management approved using Hall-B magnet for the g2p/gep expts.

  • C. Keith:

1) Working on Hall-B magnet cryostat modifications

  • J. Zhang:

1) No decision was made on where to locate the NMR coil - inside the target or outside of it?
2) Chris says it is better to keep it inside for accurate NMR measurement of the target material which sees the beam.
3) Dilutions due to NMR coil can be studied using an empty cell with NMR coil inside.
4) A subroutine to calculate the field map of Hall-B magnet was obtained from the Hall-B simulation s/w.

  • D. Crabb:

1) Trying to get the field maps for the Hall-B magnet from Oxford. They will provide a TOSCA readable input file.
2) Might get the field map files in about weeks time.

10/20/2011: Attendees: C.Keith, Ed Folts, K.Allada, J.Zhang, JP.Chen, K. Slifer, M.Huang, P. Zhu, C. Gu, M. Cummings, A. Camsonne, D. Higinbotham, J. Benesch et. al
On Phone: J.Maxwell, D. Crabb, G.Ron

  • C. Keith:

1) Target magnet was taken apart to look at the quench problem from last month

  • a) Investigating shorts to grounds that were found
  • b) Superconducting wires were destroyed
  • c) Emailed photos to Oxford Instruments
  • d) Looking into the feasibility of rewinding the coils

2) Other option is to use Hall-B magnet which has same opening in the transverse direction

  • a) Cryostat/cryogenics needs to change to fit in Hall-A
  • b) Coils are physically smaller than the current magnet
  • c) Central bore is same, opening in 0 and 90 degrees is same
  • d) Windows on scattering chamber may need to change to accommodate both g2p/gep settings.
  • e) A feasibility study is being carried out to look into using HallB magnet

3) Hall-B magnet will work for 90degree setting (good for g2p) but cannot run in 20degrees (gep). It may have enough opening in 7 degrees.
4) GEp folks are looking to see if 7 degrees is good enough to run the gep expt. - may have to take additional elastic data at 90deg to beat down systematics. Looks like it can be done (final report next week)

  • J.P. Chen:

1) Javier Gomez postponed the g2p/gep readiness review to a later date due the uncertainties in the expt. start date.
2) He also requested document discussing impact to the g2p/gep expts. if Hall-B magnet is used.
3) A feasibility study of using HallB magnet is being done by the target group. Things such as time estimates, cryostat etc. are looked into.
4) A report will be ready next week to present to the JLab management.

  • Don Crabb:

1) Process for irradiation of target material is moving forward
2) UVA will do target tests with 2.5T after the irradiation.

09/22/2011: Attendees: C.Keith, Ed Folts, K.Allada, J.Zhang, JP.Chen, T.Michalski, K. Slifer, J. Pierce, M.Huang, P. Zhu, C. Gu, J. Gomez, N. Kalantarians, A. Camsonne
On Phone: J.Maxwell, D. Crabb

  • C. Keith:

1) Magnet cool down was done on Monday(19th). Some vacuum leaks found, probably in the magnet (Indium seals?), but no major issues were observed.
2) Magnet was ramped up yesterday and it quenched at 60 Amps. This quench happened when it was sitting at a set point and not during the ramp-up time.
3) The magnet quenched again today at 8 Amps. The symptoms looks similar to what happened during SANE expt. in Hall-C.
4) Probably a section of coil became normal (non-superconducting), may be burnt.
5) Currently ran out of idea of what to look for to fix this, other than opening the magnet and looking at the coil.
6) Will keep the magnet cold today with liquid He and do some diagnostics today and tomorrow.
7) Significant delay is expected if this problem turns out to be serious one, will have to wait and see.

  • D. Crabb:

1) Will talk to magnets person at Oxford Instruments about the quench, and some ideas how to fix the magnet.

4/21/2011: Attendees: C.Keith, Ed Folts, K.Allada, J.Zhang, JP.Chen, Al Gavalya, T.Michalski, K. Slifer, J. Pierce, M.Huang, P. Zhu, M. Cummings, C. Gu et. al
On Phone: J.Maxwell

  • C. Keith:

1) Magnet cooling testing done. It wasn't a complete success.
2) Saw vacuum leaks three times.
3) Reasons for the leaks not completely understood.
4) Good: magnet works! Bad: mysterious leaks.
5) At 4k everything looked fine. Will check transfer lines for next cool down.
6) All designs for 1K refrigerator complete, parts in the machine shop. Ordering some cables.
7) Cost estimate: microwave parts, inserts, refrigerator parts etc : 50K
8) Helium for future cool down : 10K

  • JP Chen/Jixie/Ed Folts:

1) Jixie will talk with Chris to come up with a time and plan to do the field mapping.
2) It will probably involve survey groups help.
3) Ed is concerned about the local beam dump delivery time.
4) Ed talked about the power supply for the septum.
5) JP mentioned that there is money from users (Rutgers, MIT )
6) Al will need final numbers for aperture for the the local dump. Jixie promises them by early next week.

  • Don Crabb:

1) Got hold of 70GHz electronics from Duke for the 2.5 T running.
2) Irradiation: still going through the bureaucracy. Did a test of chlorine recovery, still some regulatory issues.
3) Estimates that irradiation may be finished in Aug.

  • K. Slifer:

1) Karl mentioned about the next collaboration meeting. When should we have it?
2) Please respond to Karl's email about possible collaboration dates.

4/21/2011: Attendees: C.Keith, Ed folts, K.Allada, J.Zhang, JP.Chen, Al Gavalya, T.Michalski
On Phone: J.Maxwell

  • JP Chen:

1) Reminded everyone to go through each sub-system and see if we are missing any potential issues due the target height offset (by 9cm).
2) Stressed on the importance of knowing the target field very well and mentioned about doing a field measurement in the target lab.
3) J. Zhang agreed to volunteer in setting up a simple device to measure field at few different places and cross-check with the existing map.
4) Ed mentioned that there is translation table in Hall A (originally from Hall-C) which was used earlier to do the same thing, and can be used again.
5) Mentioned about getting a short write-up ready by 28th April for the review committee.

  • C. Keith:

1) The magnet is shipped to JLab and currently somewhere in the US.
2) All electronics for the 2.5T field have not been located yet(?)
3) Duke EIO is better and will minimize the swap time between 5T and 2.5T electronics.
4) Finished modifications to the gas panel, wiring for the magnet and waiting for the valves.
5) Refrigerator design is 99.9% finished.
6) James is looking into the lifting mechanism
7) Other small items are ordered (thermometer etc..)

  • Al Gavalya:

1) Mentioned about the items bought/machined by collaborators (Rutgers, Israel group etc..)
2) There is 8 week delay for Tungsten
3) Al will need updated run list containing the sequence in which the runs at different settings are planned.
4) JP and Tim will provide the necessary run plan to Al.

Headline text